on july 15, 2019, the hyde leather company sells one of its extra tanning machines. the machine cost 72,000$ with accumulated depreciation depreciation of 65,000$ at july 15th. (the accumulated depreciation balance is after the has been brought up to date). the machine was sold for 8,000$ cash.

1)what is the book value of the machine at the time of the sale?

2) what is the gain or loss on the sale of this machine? (INDICATE IF IT IS A GAIN OR LOSS)

3)PREPARE JOURNAL ENTRY IN GOOD FORM to record the sale of this ,machine.
Answer #1

1) Book value of machine = Cost-Accumulated depreciation = 72000-65000 = 7000

2) Gain or loss on sale of machine = 8000-7000 = 1000

3) Journal entry

Date account and explanation Debit Credit
Cash 8000
Accumulated depreciation-machine 65000
Gain on sale of machine 1000
Machine 72000